# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 5.1.1. Offline Evaluation
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$49\pm.01$
.$34\pm.00$
.$54\pm.01$
.$57\pm.01$
.$57\pm.01$
.$70\pm.01$
.$\underline{.81\pm.01}$
.$\textbf{.89}\pm\textbf{.01}.$
.$\uparrow.19$
$M_{2}$
.$45\pm.00$
.$37\pm.00$
.$56\pm.02$
.$58\pm.01$
.$60\pm.01$
.$74\pm.01$
.$\underline{.77\pm.01}$
.$\textbf{.84}\pm\textbf{.01}.$
.$\uparrow.10$
$M_{3}$
.$56\pm.03$
.$37\pm.00$
.$52\pm.03$
.$57\pm.01$
.$62\pm.02$
.$64\pm.00$
.$\underline{.73\pm.01}$
.$\textbf{.76}\pm\textbf{.01}.$
.$\uparrow.12$
$M_{4}$
.$54\pm.03$
.$39\pm.00$
.55±.01plus-or-minus.55.01.$5$01
.$58\pm.03$
.$58\pm.01$
.$72\pm.01$
@@
.$\underline{.75\pm.00}$
.$\uparrow.05$
$M_{5}$
.$54\pm.02$
.$53\pm.00$
.$58\pm.02$
.$61\pm.02$
.$59\pm.02$
.$64\pm.00$
.$\textbf{.72}\pm\textbf{.01}.$
.$\underline{.67\pm.01}$
.$\uparrow.08$
$M_{6}$
.$52\pm.02$
.44±.00plus-or-minus.44.00.$4$00
.$59\pm.03$
.$60\pm.01$
.$57\pm.01$
.$67\pm.00$
.$\textbf{.75}\pm\textbf{.01}.$
.$\underline{.73\pm.02}$
.$\uparrow.08$
$M_{7}$
.$56\pm.04$
.$42\pm.00$
.$63\pm.02$
.$64\pm.01$
.$67\pm.01$
.$78\pm.01$
.$\underline{79\pm.00}$
@@
.$\uparrow.10$
$M_{8}$
.$56\pm.03$
.$54\pm.00$
.$58\pm.02$
.$57\pm.02$
.$62\pm.02$
.66±.02plus-or-minus.66.02.$6$02
.$\textbf{.80}\pm\textbf{.01}.$
.$\underline{.77\pm.01}$
.$\uparrow.14$
$M_{9}$
.$54\pm.02$
.$45\pm.00$
.$48\pm.02$
.$52\pm.02$
.$48\pm.01$
.$57\pm.01$
.$\textbf{.61}\pm\textbf{.00}.$
.$\underline{.58\pm.02}$
.$\uparrow.04$
$M_{10}$
.$56\pm.06$
.$52\pm.00$
.$57\pm.01$
.$60\pm.01$
.$60\pm.01$
.$65\pm.01$
@@
.$\uparrow.12$
$M_{11}$
.$59\pm.03$
.$48\pm.00$
.$54\pm.01$
.$60\pm.01$
.$54\pm.01$
.$63\pm.00$
.$\underline{.66\pm.00}$
.$\textbf{.68}\pm\textbf{.01}.$
.$\uparrow.05$
$M_{12}$
.$60\pm.02$
.$51\pm.00$
.55±.04plus-or-minus.55.04.$5$04
.$61\pm.02$
.$59\pm.02$
.$70\pm.01$
@@
.$\underline{.77\pm.01}$
.$\uparrow.10$
$M_{13}$
.$61\pm.01$
.44±.00plus-or-minus.44.00.$4$00
.$60\pm.02$
.$59\pm.04$
.$64\pm.01$
.$67\pm.00$
.$\textbf{.70}\pm\textbf{.00}.$
.$\underline{.68\pm.00}$
.$\uparrow.03$
$M_{14}$
.$60\pm.02$
.$52\pm.00$
.66±.01plus-or-minus.66.01.$6$01
.$68\pm.02$
.$65\pm.00$
.$65\pm.00$
.$\underline{.69\pm.00}$
.$\textbf{.70}\pm\textbf{.00}.$
.$\uparrow.05$
$M_{15}$
.$57\pm.03$
.$49\pm.00$
.$60\pm.01$
.$63\pm.02$
.$60\pm.01$
.$70\pm.01$
.$\underline{.71\pm.01}$
.$\textbf{.73}\pm\textbf{.01}.$
.$\uparrow.03$
$M_{16}$
.$53\pm.02$
.$39\pm.00$
.$52\pm.02$
.$51\pm.03$
.$61\pm.01$
.$\underline{.75\pm.01}$
.$74\pm.02$
.$\textbf{.95}\pm\textbf{.02}.$
.$\uparrow.20$
Blocks
PP-GCN
EventX
KPGNN
QSGNN
[…]
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$48\pm.00$
.11±.00plus-or-minus.11.00.$1$00
.$54\pm.01$
.$56\pm.01$
.55±.01plus-or-minus.55.01.$5$01
.$70\pm.00$
.$\underline{.81\pm.01}$
.$\textbf{.89}\pm\textbf{.01}.$
.$\uparrow.19$
$M_{2}$
.44±.02plus-or-minus.44.02.$4$02
.$12\pm.00$
.55±.01plus-or-minus.55.01.$5$01
.$57\pm.01$
.$60\pm.01$
.$74\pm.00$
.$\underline{.77\pm.01}$
.$\textbf{.84}\pm\textbf{.00}.$
.$\uparrow.10$
$M_{3}$
.55±.03plus-or-minus.55.03.$5$03
.11±.00plus-or-minus.11.00.$1$00
.55±.02plus-or-minus.55.02.$5$02
.$56\pm.02$
.$62\pm.01$
.$64\pm.00$
.$\underline{.73\pm.01}$
[…]
.$\uparrow.13$
$M_{4}$
.$54\pm.04$
.$14\pm.00$
.55±.01plus-or-minus.55.01.$5$01
.$57\pm.03$
.$60\pm.01$
.$67\pm.00$
.$\textbf{.76}\pm\textbf{.01}.$
.$\underline{.75\pm.02}$
.$\uparrow.09$
$M_{5}$
.$53\pm.02$
.$24\pm.00$
.$57\pm.01$
.$59\pm.01$
.$57\pm.01$
.$64\pm.00$
.$\textbf{.70}\pm\textbf{.01}.$
.$\underline{.67\pm.01}$
.$\uparrow.06$
$M_{6}$
.$50\pm.03$
.$15\pm.00$
.$57\pm.02$
.$59\pm.01$
.$56\pm.01$
.$57\pm.00$
.$\textbf{.75}\pm\textbf{.01}.$
.$\underline{.72\pm.01}$
.$\uparrow.18$
$M_{7}$
.55±.04plus-or-minus.55.04.$5$04
.$12\pm.00$
.$61\pm.02$
.$63\pm.01$
.$62\pm.02$
.$67\pm.00$
.$\underline{.79\pm.00}$
.$\textbf{.87}\pm\textbf{.01}.$
.$\uparrow.20$
$M_{8}$
.55±.02plus-or-minus.55.02.$5$02
.$21\pm.00$
.$57\pm.02$
.55±.02plus-or-minus.55.02.$5$02
.$61\pm.01$
.$62\pm.00$
.$\textbf{.78}\pm\textbf{.02}.$
.$\underline{.76\pm.02}$
.$\uparrow.16$
$M_{9}$
.$48\pm.03$
.$16\pm.00$
.$46\pm.02$
.$46\pm.02$
.$47\pm.01$
.$52\pm.00$
.$\textbf{.60}\pm\textbf{.01}.$
.$\underline{.57\pm.02}$
.$\uparrow.08$
$M_{10}$
.55±.04plus-or-minus.55.04.$5$04
.$19\pm.00$
.$56\pm.02$
.$58\pm.01$
.$59\pm.01$
.$60\pm.00$
.$\underline{.75\pm.02}$
.$\textbf{.75}\pm\textbf{.01}.$
.$\uparrow.15$
$M_{11}$
.$57\pm.02$
.$18\pm.00$
.$53\pm.01$
.$59\pm.02$
.$52\pm.01$
.$54\pm.00$
.$\underline{.64\pm.00}$
.$\textbf{.68}\pm\textbf{.01}.$
.$\uparrow.14$
$M_{12}$
.$58\pm.02$
.$20\pm.00$
.$56\pm.02$
.$59\pm.02$
.$62\pm.02$
.$59\pm.00$
.$\textbf{.79}\pm\textbf{.01}.$
.$\underline{.77\pm.01}$
.$\uparrow.17$
$M_{13}$
.$59\pm.02$
.$15\pm.00$
.$60\pm.02$
.$58\pm.03$
.$65\pm.01$
.$64\pm.00$
.$\textbf{.69}\pm\textbf{.00}.$
.$\underline{.67\pm.00}$
.$\uparrow.04$
$M_{14}$
.$59\pm.01$
.22±.00plus-or-minus.22.00.$2$00
.$65\pm.00$
.$67\pm.02$
.$67\pm.01$
.$65\pm.00$
.$\underline{.68\pm.00}$
.$\textbf{.70}\pm\textbf{.01}.$
.$\uparrow.03$
$M_{15}$
.55±.03plus-or-minus.55.03.$5$03
.22±.00plus-or-minus.22.00.$2$00
.$58\pm.02$
.$61\pm.00$
.$57\pm.01$
.$65\pm.00$
.$\underline{.70\pm.01}$
.$\textbf{.71}\pm\textbf{.01}.$
.$\uparrow.06$
$M_{16}$
.$52\pm.02$
.$10\pm.00$
.$50\pm.01$
.$50\pm.03$
.$59\pm.01$
.$68\pm.00$
.$\underline{.73\pm.02}$
.$\textbf{.94}\pm\textbf{.02}.$
.$\uparrow.27$
5.1.2. Online Evaluation
[…]
As depicted in Figure 4, to intuitively compare and further demonstrate the effectiveness of our proposed $\mathrm{RPLM}_{SED}$ model, we select message representations of three message blocks ($M_{2}$, $M_{9}$, and $M_{19}$) from the Events2012 dataset. These are subsequently subjected to dimensionality reduction and visualization using t-SNE. Considering the long-tail distribution characteristic of social data, our attention is focused on the top seven events by quantity. From the visualization results obtained from the $\mathrm{RPLM}_{SED}$ (Figure 4 (d), (h), and (l)), it is clearly observable that event clusters exhibit stronger cohesion. Simultaneously, the separation between distinct event clusters is also more pronounced. This reflects the importance of the proposed clustering constraint, which effectively enhances the distinguishability of message representations, thereby improving the accuracy of event detection.
Blocks
PP-GCN
EventX
KPGNN
QSGNN
[…]
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$27\pm.03$
.$01\pm.00$
.$29\pm.02$
.$30\pm.01$
.$32\pm.02$
.$60\pm.01$
.$\underline{.85\pm.01}$
.$\textbf{.94}\pm\textbf{.01}.$
.$\uparrow.34$
$M_{2}$
.$21\pm.01$
.$01\pm.00$
.$37\pm.01$
.$38\pm.02$
.$42\pm.02$
.$58\pm.01$
.$\underline{.79\pm.02}$
.$\textbf{.89}\pm\textbf{.01}.$
.$\uparrow.31$
$M_{3}$
.$38\pm.03$
.$01\pm.00$
.$39\pm.04$
.$36\pm.02$
.$47\pm.01$
.$50\pm.02$
.$\underline{.76\pm.01}$
.$\textbf{.81}\pm\textbf{.00}.$
.$\uparrow.31$
$M_{4}$
.$35\pm.05$
.$01\pm.00$
.$36\pm.04$
.$36\pm.01$
.$40\pm.03$
.$51\pm.02$
.$\textbf{.64}\pm\textbf{.01}.$
.$\underline{.64\pm.02}$
.$\uparrow.13$
$M_{5}$
.$30\pm.01$
.$03\pm.00$
.$37\pm.02$
.$36\pm.02$
.$41\pm.01$
.$48\pm.01$
.$\underline{.49\pm.03}$
.$\textbf{.49}\pm\textbf{.01}.$
.$\uparrow.01$
$M_{6}$
.$27\pm.02$
.$01\pm.00$
.$35\pm.04$
.$36\pm.01$
.$35\pm.02$
.$46\pm.02$
.$\textbf{.63}\pm\textbf{.02}.$
.$\underline{.60\pm.02}$
.$\uparrow.17$
$M_{7}$
.$38\pm.05$
.$01\pm.00$
.$37\pm.02$
.$38\pm.02$
.$43\pm.01$
.$61\pm.01$
.$\underline{79\pm.00}$
.$\textbf{.92}\pm\textbf{.01}.$
.$\uparrow.31$
$M_{8}$
.$38\pm.04$
.$01\pm.00$
.$38\pm.02$
.$35\pm.02$
.$45\pm.02$
.$47\pm.01$
.$\textbf{.75}\pm\textbf{.02}.$
.$\underline{.67\pm.01}$
.$\uparrow.28$
$M_{9}$
.$32\pm.04$
.$01\pm.00$
.$23\pm.02$
.$30\pm.04$
.$26\pm.01$
.$39\pm.01$
[…]
.$\underline{.41\pm.01}$
.$\uparrow.04$
$M_{10}$
.$37\pm.04$
.$02\pm.00$
.$38\pm.02$
.$37\pm.02$
.44±.03plus-or-minus.44.03.$4$03
.$\underline{.52\pm.01}$
.$51\pm.02$
[…]
.$\uparrow.03$
$M_{11}$
.$37\pm.04$
.$01\pm.00$
.$25\pm.02$
.$24\pm.03$
.$27\pm.01$
.$\underline{.40\pm.01}$
.$29\pm.01$
.$\textbf{.63}\pm\textbf{.02}.$
.$\uparrow.23$
$M_{12}$
.$39\pm.03$
.$02\pm.00$
.$46\pm.02$
.$40\pm.04$
.$45\pm.01$
.$\underline{.57\pm.02}$
[…]
.$48\pm.01$
.$\uparrow.20$
$M_{13}$
.$39\pm.01$
.$01\pm.00$
.$36\pm.05$
.33±.03plus-or-minus.33.03.$3$03
.$42\pm.03$
.$49\pm.04$
.$\underline{.61\pm.01}$
.$\textbf{.64}\pm\textbf{.01}.$
.$\uparrow.15$
$M_{14}$
.$39\pm.06$
.$03\pm.00$
.$50\pm.03$
.$47\pm.03$
.$50\pm.03$
.$\textbf{.52}\pm\textbf{.02}.$
.$46\pm.02$
.$\underline{.50\pm.01}$
.$\downarrow.02$
$M_{15}$
.$40\pm.06$
.$02\pm.00$
.$37\pm.02$
.$32\pm.02$
.$38\pm.02$
.$49\pm.01$
.$\underline{.51\pm.01}$
.$\textbf{.63}\pm\textbf{.01}.$
.$\uparrow.14$
$M_{16}$
.$26\pm.03$
.$01\pm.00$
.$26\pm.02$
.$25\pm.02$
.$30\pm.03$
.$\underline{.46\pm.02}$
.$38\pm.03$
.$\textbf{.97}\pm\textbf{.01}.$
.$\uparrow.51$

### v2

@@ 5.1.1. Offline Evaluation
$M_{1}$$49\pm.01$$34\pm.00$$54\pm.01$$57\pm.01$$57\pm.01$$70\pm.01$$\underline{.81\pm.01}$$\textbf{.89}\pm\textbf{.01}.$$\uparrow.19$
$M_{2}$$45\pm.00$$37\pm.00$$56\pm.02$$58\pm.01$$60\pm.01$$74\pm.01$$\underline{.77\pm.01}$$\textbf{.84}\pm\textbf{.01}.$$\uparrow.10$
$M_{3}$$56\pm.03$$37\pm.00$$52\pm.03$$57\pm.01$$62\pm.02$$64\pm.00$$\underline{.73\pm.01}$$\textbf{.76}\pm\textbf{.01}.$$\uparrow.12$
$M_{4}$$54\pm.03$$39\pm.00$$5$01 .$58\pm.03$$58\pm.01$$72\pm.01$$7$01 .$\underline{.75\pm.00}$$\uparrow.05$
$M_{5}$$54\pm.02$$53\pm.00$$58\pm.02$$61\pm.02$$59\pm.02$$64\pm.00$$\textbf{.72}\pm\textbf{.01}.$$\underline{.67\pm.01}$$\uparrow.08$
$M_{6}$$52\pm.02$$4$00 .$59\pm.03$$60\pm.01$$57\pm.01$$67\pm.00$$\textbf{.75}\pm\textbf{.01}.$$\underline{.73\pm.02}$$\uparrow.08$
$M_{7}$$56\pm.04$$42\pm.00$$63\pm.02$$64\pm.01$$67\pm.01$$78\pm.01$$\underline{79\pm.00}$$8$00 .$\uparrow.10$
$M_{8}$$56\pm.03$$54\pm.00$$58\pm.02$$57\pm.02$$62\pm.02$$6$02 .$\textbf{.80}\pm\textbf{.01}.$$\underline{.77\pm.01}$$\uparrow.14$
$M_{9}$$54\pm.02$$45\pm.00$$48\pm.02$$52\pm.02$$48\pm.01$$57\pm.01$$\textbf{.61}\pm\textbf{.00}.$$\underline{.58\pm.02}$$\uparrow.04$
@@
$M_{11}$$59\pm.03$$48\pm.00$$54\pm.01$$60\pm.01$$54\pm.01$$63\pm.00$$\underline{.66\pm.00}$$\textbf{.68}\pm\textbf{.01}.$$\uparrow.05$
@@
$M_{13}$$61\pm.01$$4$00 .$60\pm.02$$59\pm.04$$64\pm.01$$67\pm.00$$\textbf{.70}\pm\textbf{.00}.$$\underline{.68\pm.00}$$\uparrow.03$
$M_{14}$$60\pm.02$$52\pm.00$$6$01 .$68\pm.02$$65\pm.00$$65\pm.00$$\underline{.69\pm.00}$$\textbf{.70}\pm\textbf{.00}.$$\uparrow.05$
$M_{15}$$57\pm.03$$49\pm.00$$60\pm.01$$63\pm.02$$60\pm.01$$70\pm.01$$\underline{.71\pm.01}$$\textbf{.73}\pm\textbf{.01}.$$\uparrow.03$
$M_{16}$$53\pm.02$$39\pm.00$$52\pm.02$$51\pm.03$$61\pm.01$$\underline{.75\pm.01}$$74\pm.02$$\textbf{.95}\pm\textbf{.02}.$$\uparrow.20$
@@
$M_{1}$$48\pm.00$$1$00 .$54\pm.01$$56\pm.01$$5$01 .$70\pm.00$$\underline{.81\pm.01}$$\textbf{.89}\pm\textbf{.01}.$$\uparrow.19$
$M_{2}$$4$02 .$12\pm.00$$5$01 .$57\pm.01$$60\pm.01$$74\pm.00$$\underline{.77\pm.01}$$\textbf{.84}\pm\textbf{.00}.$$\uparrow.10$
$M_{3}$$5$03 .11±.00plus-or-minus.11.00.$1$00 .55±.02plus-or-minus.55.02.$5$02 .$56\pm.02$$62\pm.01$$64\pm.00$$\underline{.73\pm.01}$$7$01 .$\uparrow.13$
$M_{4}$$54\pm.04$$14\pm.00$$5$01 .$57\pm.03$$60\pm.01$$67\pm.00$$\textbf{.76}\pm\textbf{.01}.$$\underline{.75\pm.02}$$\uparrow.09$
$M_{5}$$53\pm.02$$24\pm.00$$57\pm.01$$59\pm.01$$57\pm.01$$64\pm.00$$\textbf{.70}\pm\textbf{.01}.$$\underline{.67\pm.01}$$\uparrow.06$
$M_{6}$$50\pm.03$$15\pm.00$$57\pm.02$$59\pm.01$$56\pm.01$$57\pm.00$$\textbf{.75}\pm\textbf{.01}.$$\underline{.72\pm.01}$$\uparrow.18$
$M_{7}$$5$04 .$12\pm.00$$61\pm.02$$63\pm.01$$62\pm.02$$67\pm.00$$\underline{.79\pm.00}$$\textbf{.87}\pm\textbf{.01}.$$\uparrow.20$
$M_{8}$$5$02 .$21\pm.00$$57\pm.02$$5$02 .$61\pm.01$$62\pm.00$$\textbf{.78}\pm\textbf{.02}.$$\underline{.76\pm.02}$$\uparrow.16$
$M_{9}$$48\pm.03$$16\pm.00$$46\pm.02$$46\pm.02$$47\pm.01$$52\pm.00$$\textbf{.60}\pm\textbf{.01}.$$\underline{.57\pm.02}$$\uparrow.08$
$M_{10}$$5$04 .$19\pm.00$$56\pm.02$$58\pm.01$$59\pm.01$$60\pm.00$$\underline{.75\pm.02}$$\textbf{.75}\pm\textbf{.01}.$$\uparrow.15$
$M_{11}$$57\pm.02$$18\pm.00$$53\pm.01$$59\pm.02$$52\pm.01$$54\pm.00$$\underline{.64\pm.00}$$\textbf{.68}\pm\textbf{.01}.$$\uparrow.14$
$M_{12}$$58\pm.02$$20\pm.00$$56\pm.02$$59\pm.02$$62\pm.02$$59\pm.00$$\textbf{.79}\pm\textbf{.01}.$$\underline{.77\pm.01}$$\uparrow.17$
$M_{13}$$59\pm.02$$15\pm.00$$60\pm.02$$58\pm.03$$65\pm.01$$64\pm.00$$\textbf{.69}\pm\textbf{.00}.$$\underline{.67\pm.00}$$\uparrow.04$
$M_{14}$$59\pm.01$$2$00 .$65\pm.00$$67\pm.02$$67\pm.01$$65\pm.00$$\underline{.68\pm.00}$$\textbf{.70}\pm\textbf{.01}.$$\uparrow.03$
$M_{15}$$5$03 .22±.00plus-or-minus.22.00.$2$00 .$58\pm.02$$61\pm.00$$57\pm.01$$65\pm.00$$\underline{.70\pm.01}$$\textbf{.71}\pm\textbf{.01}.$$\uparrow.06$
$M_{16}$$52\pm.02$$10\pm.00$$50\pm.01$$50\pm.03$$59\pm.01$$68\pm.00$$\underline{.73\pm.02}$$\textbf{.94}\pm\textbf{.02}.$$\uparrow.27$
5.1.2. Online Evaluation
In the online detection scenario, we utilize the data from the first week as the $M_{0}$ to train the initial model, constructing subsequent message blocks daily. It is crucial to note that for offline baselines like PP-GCN, direct adaptation to the online scenario is not feasible. We circumvent this by retraining the model using previous message blocks as the training set and predicting the current message block. Unlike offline detection, the $\mathrm{RPLM}_{SED}$ method enters the maintenance phase after predicting the message block within the window. During the maintenance phase, $\mathrm{RPLM}_{SED}$ initially updates the parameters of the PLMs section through a parameter weighting method, thereby enabling rapid adaptation to new knowledge while preserving historical knowledge. Subsequently, the model is maintained using message blocks within the current window.
@@ 5.1.2. Online Evaluation
As depicted in Figure LABEL:fig:visualization, to intuitively compare and further demonstrate the effectiveness of our proposed $\mathrm{RPLM}_{SED}$ model, we select message representations of three message blocks ($M_{2}$, $M_{9}$, and $M_{19}$) from the Events2012 dataset. These are subsequently subjected to dimensionality reduction and visualization using t-SNE. Considering the long-tail distribution characteristic of social data, our attention is focused on the top seven events of each message block by quantity. From the visualization results obtained from the $\mathrm{RPLM}_{SED}$ (Figure LABEL:fig:visualization (d), (h), and (l)), it is clearly observable that event clusters exhibit stronger cohesion. Simultaneously, the separation between distinct event clusters is also more pronounced. This reflects the importance of the proposed clustering constraint, which effectively enhances the distinguishability of message representations, thereby improving the accuracy of event detection.
@@
$M_{1}$$27\pm.03$$01\pm.00$$29\pm.02$$30\pm.01$$32\pm.02$$60\pm.01$$\underline{.85\pm.01}$$\textbf{.94}\pm\textbf{.01}.$$\uparrow.34$
$M_{2}$$21\pm.01$$01\pm.00$$37\pm.01$$38\pm.02$$42\pm.02$$58\pm.01$$\underline{.79\pm.02}$$\textbf{.89}\pm\textbf{.01}.$$\uparrow.31$
$M_{3}$$38\pm.03$$01\pm.00$$39\pm.04$$36\pm.02$$47\pm.01$$50\pm.02$$\underline{.76\pm.01}$$\textbf{.81}\pm\textbf{.00}.$$\uparrow.31$
$M_{4}$$35\pm.05$$01\pm.00$$36\pm.04$$36\pm.01$$40\pm.03$$51\pm.02$$\textbf{.64}\pm\textbf{.01}.$$\underline{.64\pm.02}$$\uparrow.13$
$M_{5}$$30\pm.01$$03\pm.00$$37\pm.02$$36\pm.02$$41\pm.01$$48\pm.01$$\underline{.49\pm.03}$$\textbf{.49}\pm\textbf{.01}.$$\uparrow.01$
$M_{6}$$27\pm.02$$01\pm.00$$35\pm.04$$36\pm.01$$35\pm.02$$46\pm.02$$\textbf{.63}\pm\textbf{.02}.$$\underline{.60\pm.02}$$\uparrow.17$
$M_{7}$$38\pm.05$$01\pm.00$$37\pm.02$$38\pm.02$$43\pm.01$$61\pm.01$$\underline{79\pm.00}$$\textbf{.92}\pm\textbf{.01}.$$\uparrow.31$
$M_{8}$$38\pm.04$$01\pm.00$$38\pm.02$$35\pm.02$$45\pm.02$$47\pm.01$$\textbf{.75}\pm\textbf{.02}.$$\underline{.67\pm.01}$$\uparrow.28$
$M_{9}$$32\pm.04$$01\pm.00$$23\pm.02$$30\pm.04$$26\pm.01$$39\pm.01$$4$01 .$\underline{.41\pm.01}$$\uparrow.04$
$M_{10}$$37\pm.04$$02\pm.00$$38\pm.02$$37\pm.02$$4$03 .$\underline{.52\pm.01}$$51\pm.02$$5$00 .$\uparrow.03$
$M_{11}$$37\pm.04$$01\pm.00$$25\pm.02$$24\pm.03$$27\pm.01$$\underline{.40\pm.01}$$29\pm.01$$\textbf{.63}\pm\textbf{.02}.$$\uparrow.23$
$M_{12}$$39\pm.03$$02\pm.00$$46\pm.02$$40\pm.04$$45\pm.01$$\underline{.57\pm.02}$$7$01 .$48\pm.01$$\uparrow.20$
$M_{13}$$39\pm.01$$01\pm.00$$36\pm.05$$3$03 .$42\pm.03$$49\pm.04$$\underline{.61\pm.01}$$\textbf{.64}\pm\textbf{.01}.$$\uparrow.15$
$M_{14}$$39\pm.06$$03\pm.00$$50\pm.03$$47\pm.03$$50\pm.03$$\textbf{.52}\pm\textbf{.02}.$$46\pm.02$$\underline{.50\pm.01}$$\downarrow.02$
$M_{15}$$40\pm.06$$02\pm.00$$37\pm.02$$32\pm.02$$38\pm.02$$49\pm.01$$\underline{.51\pm.01}$$\textbf{.63}\pm\textbf{.01}.$$\uparrow.14$
$M_{16}$$26\pm.03$$01\pm.00$$26\pm.02$$25\pm.02$$30\pm.03$$\underline{.46\pm.02}$$38\pm.03$$\textbf{.97}\pm\textbf{.01}.$$\uparrow.51$
5.2. Performance Evaluation on Low-Resource Language
Low-resource social event detection often faces the challenge of scarce training resources. Recent research introduced the CLKD (Ren et al., 2024) method, utilizing a knowledge distillation-based framework (Gupta et al., 2016) to address this challenge. Specifically, CLKD initially employs pre-trained word embedding algorithms (Mikolov et al., 2013; Pennington et al., 2014; Bojanowski et al., 2017) to generate initial embeddings of messages, followed by mapping the data of low-resource languages into the word embedding space of high-resource languages through linear (Lample et al., 2018; Artetxe et al., 2018) or non-linear (Mohiuddin et al., 2020) mapping techniques. Subsequently, a teacher model trained in a high-resource language (e.g., English) guides the student model in low-resource languages (e.g., Arabic) for learning and training. However, similar to other GNN-based methods, the backbone encoder of CLKD is still GNN. GNNs learn representations from explicit message graphs, within which some noisy edges are inevitably present. This results in the knowledge learned by the teacher model not being entirely effective and accurate, which may adversely affect the learning process of the student model. To validate our hypothesis, we conduct experiments on the low-resource Arabic-Twitter dataset. We choose TwHIN-BERT (TwHIN-BERT-large) as the backbone architecture for $\mathrm{RPLM}_{SED}$ and implement CLKD by using open-source code [^6] https://github.com/RingBDStack/CLKD.